# Robust Bayesian Model Averaging for Linear Regression Models With Heavy-Tailed Errors

### v1

@@ Abstract
In this article, our goal is to develop a method for Bayesian model averaging in linear regression models to accommodate heavier tailed error distributions than the normal distribution. Motivated by the use of the Huber loss function in presence of outliers, Park and Casella, (2008) proposed the concept of the Bayesian Huberized lasso, which has been recently developed and implemented by Kawakami and Hashimoto, (2023), with hyperbolic errors. Because the Huberized lasso cannot enforce regression coefficients to be exactly zero, we propose a fully Bayesian variable selection approach with spike and slab priors, that can address sparsity more effectively. Furthermore, while the hyperbolic distribution has heavier tails than a normal distribution, its tails are less heavy in comparison to a Cauchy distribution. Thus, we propose a regression model, with an error distribution that encompasses both hyperbolic and Student-$t$ distributions. Our model aims to capture the benefit of using Huber loss, but it can also adapt to heavier tails, and unknown levels of sparsity, as necessitated by the data. We develop an efficient Gibbs sampler with Metropolis Hastings steps for posterior computation. Through simulation studies, and analyses of the benchmark Boston housing dataset and NBA player salaries in the 2022-2023 season, we show that our method is competitive with various state-of-the-art methods.
@@ 1 Introduction
The hyperbolic distribution has heavier tails than a normal distribution, but has thinner tails than some members of the Student-$t$ family, such as the Cauchy distribution. Therefore, we propose a model in which the errors follow either a hyperbolic or a Student-$t$ distribution. With this flexible modeling strategy, we can take advantage of the Huber loss function via the hyperbolic distribution component of the error model. However, if necessary, the model can also seamlessly adapt to the heavier tails of the Student-$t$ distribution, as dictated by the data. A model with Student-$t$ errors has been developed by Gramacy and Pantaleo, (2010), who also put spike and slab priors on the regression coefficients. They used a continuous prior for the tail heaviness parameter and developed a reversible jump MCMC sampler for posterior computation. In Sections 4 and 5, we compare our method to those of Gramacy and Pantaleo, (2010) and Kawakami and Hashimoto, (2023), and demonstrate that our method is competitive with these state-of-the-art methods.
@@
Before proceeding further, the organization of the present article is outlined as follows. In Section 2, we review the hyperbolic error model. In Section 3, we introduce our proposed mixture model using hyperbolic and Student-$t$ errors. In Section 4, we carry out a simulation study to compare the performance of the proposed mixture model with some competing models. We analyze the Boston housing dataset and NBA player salaries in Section 5. Concluding remarks and possible directions for future work are stated in Section 6. Additionally, for clarity on parametrization, the probability density functions of some distributions that are used throughout the article, are provided in an Appendix.
@@ 2.1 Review of the Hyperbolic Distribution
The hyperbolic density in (4) has a heavier tail than the normal density, but is lighter-tailed than a Cauchy distribution. This is illustrated in Figure 1, where the cumulative distribution functions of hyperbolic distributions (for different choices of $\eta$ and $\rho^{2}=1$) are compared with those for standard normal and Student $t$-distributions with one degree of freedom (that is, the standard Cauchy distribution). Furthermore, we note that the $\mathrm{Hyperbolic}(\eta=50,\rho^{2}=1)$ distribution is almost identical to the standard normal distribution, which provides useful information for specifying a prior for $\eta$ in (19), in Section 3.1.
@@ 2.2 Model Formulation
We close our deliberation on the hyperbolic error model through a summary of its significance and pitfalls. To deal with heavy-tailed data, a plausible contender is the Student-$t$ error model. However, the simulation study in Section 4.1 reveals the usefulness of the hyperbolic error model, when the true data generating model is hyperbolic. Our simulation study suggests that the hyperbolic model can have superior performance in estimation of regression coefficients in such a case. These results provide another motivation for using the hyperbolic error model, in addition to its relationship with the Huber loss. On the other hand, the Student-$t$ distribution can achieve a higher degree of tail heaviness compared to the hyperbolic distribution. Accordingly, both the error models have some benefits and drawbacks, entailing a probe for a more befitting error model.
@@ 3 Mixture of Hyperbolic and Student-$t$ Error Models
It has been discussed in Section 2 that despite having heavier tails than a normal distribution, the hyperbolic distribution has thinner tails than some members of the family of Student-$t$ distributions. However, there are situations where a hyperbolic error model provides better estimation than a Student-$t$ error model, like the one shown in Section 4.1. This suggests that it would be appealing to have a model that accounts for the characteristics of both distributions. Inspired by this idea, we propose a linear regression model, where the errors follow a mixture of hyperbolic and Student-$t$ distributions. We incorporate variable selection uncertainty in our model, via spike and slab priors on the regression coefficients.
@@ 3.2 Posterior Computation
where $A_{\bm{\gamma}}=X_{\bm{\gamma}}^{\top}D^{-1}X_{\bm{\gamma}}+\dfrac{1}{\tau^{2}\rho^{2}}I_{p_{\bm{\gamma}}}$, $p_{\bm{\gamma}}=\sum_{j=1}^{p}\bm{\gamma}_{j}$, and ${\epsilon}_{i}=y_{i}-\bm{x}^{(\bm{\gamma})\top}_{i}\bm{\beta}_{\bm{\gamma}}$, for $i=1,2,\dots,n$. With a current model state at $\bm{\gamma}$, the acceptance probability of a proposed model $\bm{\gamma}^{*}$ in the Metropolis-Hastings step is computed as
@@ 4 Simulation Study
$y_{i}=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}+\dots+\beta_{p}x_{ip}+\epsilon_{i},\quad i=1,2,\dots,p,$ (30)
@@ 4.2 Comparison of the HTEM With Existing Methods
This section ventures to explore the advantages of the HTEM vis-à-vis some prevailing candidate methods. The Huberized lasso (Park and Casella, (2008)), being a stemming entity for our proposed method, asks for consideration in our comparison. The Bayesian Huberized lasso (HBL) of Kawakami and Hashimoto, (2023) uses an approximate Gibbs sampler. Kawakami and Hashimoto, (2023) put a gamma prior on the shape parameter $\eta$. This gives rise to a full conditional distribution for $\eta$, that is not available in closed form. Thus, Kawakami and Hashimoto, (2023) approximate the full conditional of $\eta$ by a gamma distribution. However, in our comparison, we refrain from using such approximation and replace the same with our own discrete uniform prior on the grid in (19). We also use the inverse gamma prior in (10) on the scale parameter $\rho^{2}$ instead of an improper prior proposed by Kawakami and Hashimoto, (2023). Such an approach is deployed to ensure a parity among the compared methods. The HBL uses Laplace priors on the regression coefficients corresponding to the lasso. With our implementation of the HBL, the comparison of the HBL with the HTEM boils down to comparing 1) Laplace versus spike and slab priors on regression coefficients, and 2) hyperbolic errors in the HBL, versus a mixture of hyperbolic and Student-$t$ errors in the HTEM. In addition to the HBL, we consider the normal and the Student-$t$ error models proposed by Gramacy and Pantaleo, (2010), which perform variable selection using spike and slab priors, through reversible jump MCMC algorithms. Both the techniques by Gramacy and Pantaleo, (2010) are implemented using the monomvn package in R, and we denote them by MONON and MONOT for the normal and Student-$t$ models respectively.
@@ 4.3 Detection of True Error Density
A pertinent query is exploring the ability of our proposed method to detect the true error model. We recall that in the proposed HTEM, an indicator $\alpha$ is introduced to switch between hyperbolic and Student-$t$ densities. In particular, the hyperbolic density is chosen for $\alpha=0$, while $\alpha=1$ denotes the Student-$t$ density. Moreover, the normal distribution can be captured under both $\alpha=0$ and $\alpha=1$, for large values of $\eta$.
We study the proportion of times a hyperbolic density has been chosen by the HTEM, which is the Monte Carlo estimate of the posterior probability of a hyperbolic error model, $P(\alpha=0|\bm{Y})$. For all the scenarios, the corresponding MCMC sample proportion is used as an estimate. The results are displayed in Figure 8. Under Scenarios (I) and (IV), when the true errors are generated from a hyperbolic distribution, the posterior probability of a hyperbolic model, $P(\alpha=0|\bm{Y})$, is estimated to be much higher than the corresponding prior probability 0.5. Similarly, under Scenarios (III) and (IV), which correspond to true Student-$t$ errors, the posterior probability of the Student-$t$ error model, given by $1-P(\alpha=0|\bm{Y})$, is usually estimated to be higher than 0.5, the prior probability. However, the (estimated) posterior probability of the correct error model, is higher under true hyperbolic errors in Scenarios (I) and (IV), than those under Scenarios (III) and (VI). As far as the true normal errors in Scenarios (II) and (V) are concerned, the estimate of the posterior probability of a hyperbolic model is close to or slightly above 0.5. The result is sensible since normal errors can arise from large values of $\eta$ from either component of the mixture distribution. The usefulness of the proposed HTEM, as discussed in Section 3.2, thus gets yet another empirical justification.
@@ 5.1 Boston Housing Dataset
Figure 10: Median absolute deviations (top panel), empirical coverage probabilities (middle panel) and median widths (bottom panel) of 90% prediction intervals for the HTEM and the other studied competing methods for the Boston Housing dataset. In the middle panel, the diamonds indicate the mean empirical coverages and the dashed line marks a coverage of 90%. The boxplots and means are based on 100 replicates.
@@ 5.2 NBA Player Salaries Dataset
Figure 12: Median absolute deviations (top panel), empirical coverage probabilities (middle panel) and median widths (bottom panel) of 90% prediction intervals for the HTEM and the other studied competing methods for the NBA Player Salaries dataset. In the middle panel, the diamonds indicate the mean empirical coverages and the dashed line marks a coverage of 90%. The boxplots and means are based on 100 replicates.

### v2

@@ Abstract
Our goal is to develop a Bayesian model averaging technique in linear regression models that accommodates heavier tailed error densities than the normal distribution. Motivated by the use of the Huber loss function in the presence of outliers, the Bayesian Huberized lasso with hyperbolic errors has been proposed and recently implemented in the literature (Park and Casella, (2008); Kawakami and Hashimoto, (2023)). Since the Huberized lasso cannot enforce regression coefficients to be exactly zero, we propose a fully Bayesian variable selection approach with spike and slab priors to address sparsity more effectively. The shapes of the hyperbolic and the Student-$t$ density functions are different. Furthermore, the tails of a hyperbolic distribution are less heavy compared to those of a Cauchy distribution. Thus, we propose a flexible regression model with an error distribution encompassing both the hyperbolic and the Student-$t$ family of distributions, along with an unknown tail heaviness parameter, that is estimated based on the data. It is known that the limiting form of both the hyperbolic and the Student-$t$ distributions is a normal distribution. We develop an efficient Gibbs sampler with Metropolis Hastings steps for posterior computation. Through simulation studies and analyses of real datasets, we show that our method is competitive with various state-of-the-art methods.
@@ 1 Introduction
The shapes of the hyperbolic and the Student-$t$ density functions are distinct. As a result, some heavy-tailed members of the hyperbolic family of distributions have the potential to generate extreme observations more frequently than some heavy-tailed members of the Student-$t$ family, such as the Cauchy distribution. However, the Cauchy distribution has heavier tails than any hyperbolic density. This implies that, while the Cauchy distribution may produce extreme observations much less frequently, they could be much more extreme in magnitude than those produced under the hyperbolic distribution. Moreover, the limiting form of both the hyperbolic and the Student-$t$ distributions is a normal distribution. Therefore, we propose a model in which the errors follow either a hyperbolic or a Student-$t$ distribution. With this flexible modeling strategy, we can take advantage of the Huber loss function via the hyperbolic distribution component of the error model. However, if necessary, the model can also seamlessly adapt to the heavier tails of the Student-$t$ distribution or to the thinner tails of the normal distribution, as dictated by the data. A model with Student-$t$ errors has been developed by Gramacy and Pantaleo, (2010), who also put spike and slab priors on the regression coefficients. They used a continuous prior for the tail heaviness parameter and developed a reversible jump MCMC sampler for posterior computation. In Sections 4 and 5, we compare our method to those of Gramacy and Pantaleo, (2010) and Kawakami and Hashimoto, (2023), and demonstrate that our method is competitive with these state-of-the-art methods.
@@
Before proceeding further, the organization of the present article is outlined as follows. In Section 2, we review the hyperbolic error model. In Section 3, we introduce our proposed mixture model using hyperbolic and Student-$t$ errors. In Section 4, we carry out a simulation study to compare the performance of the proposed mixture model with some competing models. We analyze the Boston housing dataset and NBA player salaries in Section 5. Concluding remarks and possible directions for future work are stated in Section 6. Additionally, for clarity on parametrization, the probability density functions of some distributions that are used throughout the article are provided in an Appendix.
@@ 2.1 Review of the Hyperbolic Distribution
The hyperbolic density in (4) has a heavier tail than the normal density, but is lighter-tailed than a Cauchy distribution. This is illustrated in Figure 1, where the cumulative distribution functions of hyperbolic distributions (for different choices of $\eta$ and $\rho^{2}=1$) are compared with those for standard normal and Student $t$-distributions with one degree of freedom (that is, the standard Cauchy distribution). Furthermore, we note that the $\mathrm{Hyperbolic}(\eta=50,\rho^{2}=1)$ distribution is almost identical to the standard normal distribution, which provides useful information for specifying a prior for $\eta$ in (19), in Section 3.1. Additionally, the hyperbolic distribution approaches a Laplace distribution as the shape or tail heaviness parameter $\eta$ tends to zero (Barndorff-Nielsen,, 1977).
@@ 2.2 Model Formulation
We close our deliberation on the hyperbolic error model through a summary of its significance and pitfalls. To deal with heavy-tailed data, a plausible contender is the Student-$t$ error model. However, the simulation study in Section 4.1 reveals the usefulness of the hyperbolic error model, when the true data generating model is hyperbolic. Our simulation study suggests that the hyperbolic model can have superior performance in estimation of regression coefficients in such a case. These results provide another motivation for using the hyperbolic error model, in addition to its relationship with the Huber loss. The convergence of the hyperbolic density to a Laplace density for small values of $\eta$ (Barndorff-Nielsen, (1977)) provides yet another motivation for including the former distribution. On the other hand, the Student-$t$ distribution can achieve a higher degree of tail heaviness compared to the hyperbolic distribution. Besides, the two distributions have contrasting tail structures. While very large observations can be sampled from the hyperbolic density more frequently than the Student-$t$ density, the Student-$t$ distribution has the potential to generate values that are even more extreme than the hyperbolic distribution. At the same time, a behavior similar to a normal distribution can be achieved by both the distributions for large values of the tail heaviness parameter. Accordingly, both the hyperbolic and the Student-$t$ error models have some benefits and drawbacks, entailing a probe for a more befitting error model.
@@ 3 Mixture of Hyperbolic and Student-$t$ Error Models
It has been discussed in Section 2 that despite having heavier tails than a normal distribution, the hyperbolic distribution has thinner tails than some members of the family of Student-$t$ distributions. However, there are situations where a hyperbolic error model provides better estimation than a Student-$t$ error model, like the one shown in Section 4.1. The two distributions also differ in shapes, as mentioned in Section 2. This suggests that it would be appealing to have a model that accounts for the characteristics of both distributions. Inspired by this idea, we propose a linear regression model, where the errors follow a mixture of hyperbolic and Student-$t$ distributions. We incorporate variable selection uncertainty in our model, via spike and slab priors on the regression coefficients.
@@ 3.2 Posterior Computation
where $A_{\bm{\gamma}}=X_{\bm{\gamma}}^{\top}\Sigma^{-1}X_{\bm{\gamma}}+\dfrac{1}{\tau^{2}\rho^{2}}I_{p_{\bm{\gamma}}}$, $p_{\bm{\gamma}}=\sum_{j=1}^{p}\bm{\gamma}_{j}$, and ${\epsilon}_{i}=y_{i}-\bm{x}^{(\bm{\gamma})\top}_{i}\bm{\beta}_{\bm{\gamma}}$, for $i=1,2,\dots,n$. With a current model state at $\bm{\gamma}$, the acceptance probability of a proposed model $\bm{\gamma}^{*}$ in the Metropolis-Hastings step is computed as
@@ 4 Simulation Study
$y_{i}=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}+\dots+\beta_{p}x_{ip}+\epsilon_{i},\quad i=1,2,\dots,n,$ (30)
@@ 4.2 Comparison of the HTEM With Existing Methods
This section ventures to explore the advantages of the HTEM vis-à-vis some prevailing candidate methods. The Huberized lasso (Park and Casella, (2008)), being a stemming entity for our proposed method, asks for consideration in our comparison. The Bayesian Huberized lasso (HBL) of Kawakami and Hashimoto, (2023) uses an approximate Gibbs sampler. Kawakami and Hashimoto, (2023) put a gamma prior on the shape parameter $\eta$. This gives rise to a full conditional distribution for $\eta$, that is not available in closed form. Thus, Kawakami and Hashimoto, (2023) approximate the full conditional of $\eta$ by a gamma distribution. However, in our comparison, we refrain from using such approximation and replace the same with our own discrete uniform prior on the grid in (19). We also use the inverse gamma prior in (10) on the scale parameter $\rho^{2}$ instead of an improper prior proposed by Kawakami and Hashimoto, (2023). Such an approach is deployed to ensure a parity among the compared methods. The HBL uses Laplace priors on the regression coefficients corresponding to the lasso. With our implementation of the HBL, the comparison of the HBL with the HTEM boils down to comparing 1) Laplace versus spike and slab priors on regression coefficients, and 2) hyperbolic errors in the HBL, versus a mixture of hyperbolic and Student-$t$ errors in the HTEM. In addition to the HBL, we consider the normal and the Student-$t$ error models proposed by Gramacy and Pantaleo, (2010), which perform variable selection using spike and slab priors, through reversible jump MCMC algorithms. Both the techniques by Gramacy and Pantaleo, (2010) are implemented using the monomvn package (Gramacy, (2024)) in R, and we denote them by MONON and MONOT for the normal and Student-$t$ models respectively.
@@
We also assess the efficacy of the HTEM in variable selection by comparing the same with the competing methods involving spike and slab priors, i.e., the MONOT and the MONON. For every $j=1,2,...,p$, we look at the inclusion indicator $\gamma_{j}$, which is estimated using the marginal posterior inclusion probability $P(\gamma_{j}=1|\bm{Y})$ as
$\hat{\gamma}_{j}=1\iff P(\gamma_{j}=1|\bm{Y})\geq\lambda,$ (31)
where $\lambda\in(0,1)$ is a suitable threshold. Since the marginal posterior inclusion probabilities are analytically intractable, we use the corresponding MCMC sample proportions as their estimates. The next question relates to the choice of the threshold $\lambda$. The guiding instrument in our case is the marginal Bayes factor (Kass and Raftery, (1995); Ghosh and Ghattas, (2015)) for $\gamma_{j}=1$ versus $\gamma_{j}=0$, defined as
$\mathrm{BF}_{j}=\dfrac{P(\gamma_{j}=1|\bm{Y})/P(\gamma_{j}=0|\bm{Y})}{P(\gamma_{j}=1)/P(\gamma_{j}=0)},$ (32)
which is the ratio of the posterior odds to the prior odds of $\gamma_{j}=1$. Aligning with the suggestions by Kass and Raftery, (1995), a marginal Bayes factor exceeding 3.2 would indicate substantial evidence in favor of including the $j$th covariate in the model. In this context, a possible option is to employ the median probability model (MPM) with $\lambda=0.5$, as proposed by Barbieri and Berger, (2004). This corresponds to a marginal Bayes factor threshold of 1 under a uniform prior on the inclusion probability $\tilde{\pi}$. However, such an approach appears to provide a bit stringent condition of variable selection for the sparse framework with a $\mathrm{Beta}(1,\sqrt{p})$ prior on $\tilde{\pi}$, considered in our study, because using the MPM with our choice of a $\mathrm{Beta}(1,\sqrt{p})$ prior leads to a very high threshold of 10 for the marginal Bayes factor. Therefore, while retaining the cutoff for marginal Bayes factor at 3.2, we use the prior under our sparse framework, leading to a threshold of $\lambda=0.24$ for the marginal posterior inclusion probabilities. Using the subsequent estimates $\hat{\gamma}_{j}$s, we compute the true positive rates (TPRs) and true negative rates (TNRs) against the true inclusion indicators, in order to assess the variable selection performances of the methods. It is to be noted that TPR denotes the proportion of signals correctly captured in a model, among all the true signals. Similarly, TNR refers to the proportion of correctly dropped noise variables among all the true noise variables. As such, both TPR and TNR should be equal to 1 in an ideal case of perfect detection of signals and noise variables. For our study, the TPRs and TNRs obtained by the HTEM, the MONOT and the MONON are summarized in Table 3, based on the 100 simulated replications. The TPR values in Table 3 are very close to 1 and are nearly comparable for all the methods under each of the six scenarios. On the other hand, the HTEM shows a significant gain in TNR. On a closer look, under scenarios corresponding to the mixed signals (Scenarios (IV), (V) and (VI)), the gain in TNR by the HTEM appears to be even more notable. In other words, the proposed HTEM shines in identifying the noise variables and appears to be slightly less effective than the other methods in identifying the signals, thereby corroborating the observations based on RMSEs of estimated regression coefficients.
Table 3: TPRs (TNRs) of all the regression coefficients for the HTEM and the other studied competing methods under different scenarios (using a threshold $\lambda$ = 0.24 for marginal inclusion probabilities). The values are averaged over 100 replications.
HTEM MONOT MONON
Scenario (I) (Hyperbolic) 1.000 (0.995) 1.000 (0.680) 1.000 (0.665)
Scenario (II) (Normal) 1.000 (0.998) 1.000 (0.917) 1.000 (0.916)
Scenario (III) (Student-t) 1.000 (0.990) 1.000 (0.933) 0.994 (0.715)
Scenario (IV) (Hyperbolic) 0.818 (0.993) 0.985 (0.208) 0.980 (0.237)
Scenario (V) (Normal) 0.953 (0.996) 1.000 (0.771) 1.000 (0.771)
Scenario (VI) (Student-t) 0.958 (0.988) 1.000 (0.756) 0.995 (0.437)
@@ 4.3 Detection of True Error Density
A pertinent query is exploring the ability of our proposed method to detect the true error model. We recall that in the proposed HTEM, an indicator $\alpha$ is introduced to switch between hyperbolic and Student-$t$ densities. In particular, the hyperbolic density is chosen for $\alpha=0$, while $\alpha=1$ denotes the Student-$t$ density. Moreover, the normal distribution can be captured under both $\alpha=0$ and $\alpha=1$, for large values of $\eta$. Accordingly, it is reasonable to look at the posterior probability $P(\alpha=0|\bm{Y})$ of a hyperbolic error model. Considering the behavior of the densities with respect to $\alpha$, we would typically expect $P(\alpha=0|\bm{Y})$ to be close to 1 and 0 respectively, depending on whether the true errors are generated from hyperbolic or Student-$t$ distributions. For true normal errors, since both $\alpha=0$ and $\alpha=1$ can lead to a normal density under large $\eta$, $P(\alpha=0|\bm{Y})$ is likely to assume a value of about 0.5.
In the present article, we study the proportion of times a hyperbolic density has been chosen by the HTEM, which is the Monte Carlo estimate of $P(\alpha=0|\bm{Y})$. For each of the scenarios, the corresponding MCMC sample proportion is used as an estimate. The results are displayed in Figure 8. Under Scenarios (I) and (IV), when the true errors are generated from a hyperbolic distribution, the posterior probability of a hyperbolic model, $P(\alpha=0|\bm{Y})$, is estimated to be much higher than the corresponding prior probability 0.5. Similarly, under Scenarios (III) and (IV), which correspond to true Student-$t$ errors, the posterior probability of the Student-$t$ error model, given by $1-P(\alpha=0|\bm{Y})$, is usually estimated to be higher than 0.5, the prior probability. However, the (estimated) posterior probability of the correct error model, is higher under true hyperbolic errors in Scenarios (I) and (IV), than those under Scenarios (III) and (VI). As far as the true normal errors in Scenarios (II) and (V) are concerned, the estimate of the posterior probability of a hyperbolic model is close to or slightly above 0.5. In other words, all the scenarios yield sensible results, which further empirically justifies the usefulness of the proposed HTEM, as discussed in Section 3.2.
@@ 5.1 Boston Housing Dataset
Figure 10: Relative MeADs (top panel), empirical coverage probabilities (middle panel) and median widths (bottom panel) of 90% prediction intervals for the HTEM and the other studied competing methods for the Boston Housing dataset. In the top panel, relative MeAD of any method is its MeAD relative to that of the method having the smallest MeAD. In the middle panel, the diamonds indicate the mean empirical coverages and the dashed line marks a coverage of 90%. The boxplots and means are based on 100 replicates.
@@ 5.2 NBA Player Salaries Dataset
Figure 12: Relative MeADs (top panel), empirical coverage probabilities (middle panel) and median widths (bottom panel) of 90% prediction intervals for the HTEM and the other studied competing methods for the NBA Player Salaries dataset. In the top panel, relative MeAD of any method is its MeAD relative to that of the method having the smallest MeAD. In the middle panel, the diamonds indicate the mean empirical coverages and the dashed line marks a coverage of 90%. The boxplots and means are based on 100 replicates.